OpenStudy (anonymous):

1,550= −x2 + 105x − 1050

OpenStudy (anonymous):

Im stuck at 2600=x^2+105x

OpenStudy (anonymous):

make it ax^2 + bx + c =0

OpenStudy (anonymous):

x² - 105x + 1050 + 1,550 = 0

OpenStudy (anonymous):

thats the answer? ^

OpenStudy (anonymous):

Apply quadratic equation to solve for x !!!

OpenStudy (anonymous):

ohhh okay thanks